# On the complexity of constrained reconfiguration and motion planning

### v1

@@ 3.2 Hardness results
The procedures that optimally solve the $k$-Compatible Ordering problem and that were presented in Section 3.1 (proofs of Lemmas 12) exploited acyclicity within the input. In this section, we restrict the presence of cycles to represent cyclical constraints that naturally arise in the context of the RAMP problems. We prove below that $k$-Compatible Ordering is hard under each of those restrictions.
@@
Membership in $\mathsf{NP}$ follows from the fact that a labeled ordering $(\mathcal{S},\mathcal{L})$ serves as a certificate; verifying all constraints in $\mathcal{O}(n+m)$ time is straightforward (Observation 5).
@@
Let $c_{i}^{a}$ be a clause vertex labeled 1, and let $x_{j}$ be the associated variable. If $c_{i}^{a}$ is a positive clause vertex, given the arc $c_{i}^{a}x_{j}$ in $A_{1}$, $x_{j}$ must come before $c_{i}^{a}$ (Observation 1). Moreover, since we can now assume that $x_{j}$ comes before $c_{i}^{a}$, $x_{j}$ cannot be assigned label 2 due to the arc $x_{j}c^{a}_{i}$ in $A_{2}$. Otherwise, if $c_{i}^{a}$ is a negative clause vertex, given the arc $x_{j}c^{a}_{i}$ in $B_{1}$, $x_{j}$ must come after $c_{i}^{a}$ (Observation 1). Moreover, since we can now assume that $x_{j}$ comes after $c_{i}^{a}$, $x_{j}$ cannot be assigned label 1 due to the arc $c^{a}_{i}x_{j}$ in $B_{1}$. ∎
Let $C^{\ell=1}$ be the set of clause vertices that are assigned label 1. It follows from Observation 8 that $C^{\ell=1}$ cannot include negative and positive clause vertices with the same associated variable, as this would imply that the associated variable is included multiple times in the ordering, and is assigned two different labels. We partition $C^{\ell=1}$ into two sets of clause vertices, $C^{+,\ell=1}$ and $C^{-,\ell=1}$, based on polarity, i.e., $C^{+,\ell=1}$ (resp. $C^{-,\ell=1}$) is the set of positive (resp. negative) clause vertices in $C^{l=1}$. Let $X^{+,\ell=1}$ (resp. $X^{-,\ell=1}$) be the set of variable vertices associated with the clause vertices in $C^{+,\ell=1}$ (resp. $C^{-,\ell=1}$). Note that $|X^{+,\ell=1}|\leq|C^{+,\ell=1}|$ and $|X^{-,\ell=1}|\leq|C^{-,\ell=1}|$, since multiple clause vertices may be associated with the same variable vertex, and that $X^{+,\ell=1}\cap X^{-,\ell=1}=\emptyset$, which follows from Observation 8.
We observe that $X\setminus\left(X^{+,\ell=1}\bigcup X^{-,\ell=1}\right)$ need not be empty. We claim that setting each variable in $X^{+,\ell=1}$ to true and setting each variable in $X^{-,\ell=1}$ to false satisfies $\varphi$; the variables in $X\setminus\left(X^{+,\ell=1}\bigcup X^{-,\ell=1}\right)$ are don’t-care variables that can be set to either true or false. Suppose otherwise: there exists a clause $C_{i}$ in $\varphi$, that is not satisfied by the proposed truth value assignment. By Observation 3, a clause vertex $c_{i}^{a}$ of $C_{1}$, whose associated variable is $x_{j}$, is included in the ordering with label 1. Furthermore, by Observation 8, if $c^{a}_{i}$ is a positive (resp. negative) clause vertex, $x_{j}\in X^{+,\ell=1}$ (resp. $x_{j}\in X^{-,\ell=1}$), and is assigned the truth assignment true (resp. false) by our construction, which satisfies $c_{i}^{1}$, and therefore $C_{i}$, a contradiction. This concludes the proof of correctness of the reduction.
@@
The construction covered in Theorem 2 goes against the spirit of the problem, because it relies on the existence of directed digons in $B_{1}$ and directed triangles in $A_{2}$: directed cycles trivially prevent the existence of a solution where the cycle’s vertices are all assigned the same label (Observation 3). One might therefore wonder whether forbidding directed cycles within each of the directed graphs makes the problem easier to solve. Our next result shows that the answer is negative.
@@
    Arcs involving the global vertex $g$: For each dummy clause vertex $d_{i}$, add the arc $d_{i}g$ to $A_{2}$ and the arc $gd_{i}$ to $B_{2}$. For each dummy negative clause vertex $d_{i}^{a}$, add the arc $d_{i}^{a}g$ to $A_{1}$ and the arc $gd_{i}^{a}$ to $B_{1}$. These arcs guarantee dummy vertices obtain particular labels, as explained in Observation 4, and prevent cycles involving $g$.
@@
($\Leftarrow$) The proof for the backward direction of Theorem 2 relied on the fact that one of $c_{i}^{1},c_{i}^{2}$ and $c_{i}^{3}$ is assigned label 1 for each clause $C_{i}$, as well as Observation 8. We show that the two observations still hold under the new construction, which is sufficient to prove the existence of a satisfying assignment for $\varphi$.
To prove that, for each clause $C_{i}$, one of $c_{i}^{1},c_{i}^{2},c_{i}^{3}$ is assigned label 1 under the new construction, suppose, for the sake of contradiction, that there exists a clause $C_{i}$ whose clause vertices are all assigned label 2 in the ordering, which implies that its non-dummy clause vertices are in the following relative order: $c_{i}^{3},c_{i}^{2},c_{i}^{1}$. Since $d_{i}$ is assigned label 1 (Observation 4), it will have to come after $c_{i}^{1}$, and therefore, after $c_{i}^{3}$, which violates a constraint on $c_{i}^{3}$ because of the arc $c_{i}^{3}d_{i}$ in $A_{1}$ (Observation 1). As for Observation 8, the previous proof is still applicable in the new construction, because the proof relies on the arcs constructed between variable vertices and non-dummy clause vertices, which were not modified.
@@
In the construction of Theorem 3, cycles no longer exist in individual directed graphs, but rather, in the union of the directed graphs. By Lemma 2, we already know that the problem admits a polynomial-time solution when the union of $\mathcal{G}$ is acyclic (and actually even when there exists a single $i$ for which $A_{i}\cup B_{i}$ is acyclic). So the natural next step would be looking for intermediate instances. Next, we prove that the problem remains $\mathsf{NP}$-complete even on instances whose union is acyclic after a simple transformation of the instance. Given a directed graph $G=(V,E)$, we define the reversed graph of $G$ as the graph $(V,\{v_{j}v_{i}$ $|$ $v_{i}v_{j}\in E\})$: we will reverse the $B$ graphs.
@@
($\Leftarrow$) Let $(\mathcal{S},\mathcal{L})$ be the valid labeled ordering that solves the constructed $3$-Compatible Ordering instance. In addition to the vertex $t_{1}$ coming before the vertex $t_{2}$ in $\mathcal{S}$, since the arc $t_{2}t_{1}$ exists in all $A$ graphs (Observation 2), $\mathcal{S}$ has the following characteristics:
Observation 9.
@@
Observation 9 is due to $t_{1}$ coming before $t_{2}$, and the arcs $C_{i}t_{1}$ in the $A$ graphs, $t_{2}C_{i}$ in the $B$ graphs for each clause vertex $C_{i}$. As for Observation 10, for each variable vertex $x_{j}$, the arc $t_{1}x_{j}$ in $B_{1}$ means that $x_{j}$ must come before $t_{1}$ if $x_{j}$ is labeled 1 (Observation 1), the arcs $x_{j}t_{1}$ and $t_{1}x_{j}$ in $A_{2}$ and $B_{2}$ respectively mean that $x_{j}$ is never labeled 2 (Observation 4), and the arc $x_{j}t_{2}$ in $A_{3}$ means that $x_{j}$ must come after $t_{2}$ if $x_{j}$ is labeled 3 (Observation 1).
In order to derive a satisfying assignment, we focus on the clause vertices, as well as the labels that they are assigned. Let $\ell_{1},\ldots,\ell_{p}$ be the labels assigned to $C_{1},\ldots,C_{p}$, respectively. Consider the set of literals $C^{sat}=\{c^{\ell_{1}}_{1},\ldots,c^{\ell_{p}}_{p}\}$. We will partition those literals into two sets $C^{T}$ and $C^{F}$, based on whether the arc between $c_{i}^{\ell_{i}}$ and its associated variable, in one direction or the other, is in an $A$ graph or a $B$ graph, respectively. We then define $X^{T}$ (resp. $X^{F}$) as the set of variables in $\varphi$ with at least one literal in $C^{T}$ (resp. $C^{F}$). We observe that a variable vertex $x$ cannot belong to both $X^{T}$ and $X^{F}$, i.e., $X^{T}\cap X^{F}=\emptyset$, because an arc from a clause vertex in $C^{sat}$ to $x$ in an $A$ graph implies that $x$ comes before $t_{1}$ in $\mathcal{S}$, and an arc from the same clause vertex in $C^{sat}$ to $x$ in a $B$ graph implies that $x$ comes after $t_{2}$ in $\mathcal{S}$ (Observations 110), a contradiction, since we already established that $t_{1}$ comes before $t_{2}$ in every valid labeled ordering.
@@ 5 Finding $k$-compatible set arrangements
The proof of Observation 12 is identical to that of Observation 11 because vertex additions and vertex removals are treated identically as far as constraint satisfaction is concerned, i.e., the removal of a vertex $v$ from $V_{i}$ is legal if and only if the addition of a vertex $v$ to $V_{i}\setminus\{v\}$ is legal.
@@
Analyzing whether the vertex removal we propose is legal will omit any mention of the $A$ graphs, because we have already established that a vertex removal cannot violate the removed vertex’s sink constraint (Observation 11):
@@
If $V_{\text{s}}=V_{\text{t}}$, there is nothing to do. Otherwise, we can refer to Observation 12 to conclude the existence of at least one vertex removal in $\mathbb{V}$. ∎
Observation 15.
@@
Suppose otherwise, for the sake of contradiction. If the removal of $v^{e}$ is the last addition/removal operation involving any of $v^{e}$ and $u^{e}$ in $\mathbb{V}$, we can find a shorter sequence of vertex sets that is a solution to the instance. If $\mathbb{V}=V_{1},\ldots,V_{j_{1}},V_{j_{1}+1},V_{j_{1}+2},\ldots,V_{q}$ is a solution to the instance, so is $\mathbb{V}^{\prime}=V_{1},\ldots,V_{j_{1}},V_{j_{1}+2}\cup\{v^{e}\},\ldots,V_{q}\cup\{v^{e}\}$. Undoing the removal of $v^{e}$ can only cause sink constraint violations for the vertices that are subsequently added or removed (Observation 13), and the only vertex whose sink constraint involves $v^{e}$ is $u^{e}$. A similar argument can be used to shorten the vertex set sequence that is a solution to the instance if the removal of $v^{e}$ is followed by the addition of $v^{e}$ to some vertex set $V_{j_{3}}$, where $j_{3}>j_{1}$. ∎
In what follows, we assume that $\mathbb{V}$ is a nontrivial shortest solution whose structure adheres to that of the solution in the statement of Observation 15.
Observation 16.
@@
We modify $\mathbb{V}$ to obtain $\mathbb{V}^{\prime}$ by moving $v^{e}$ directly after the removal operation of $u^{e}$, which can be thought of as the undoing of a vertex addition of $v^{e}$, followed by a vertex addition of $v^{e}$, both to a sequence of vertex sets. This modification can only violate sink constraints (Observation 13), and the only vertex that is potentially affected by the earlier inclusion of $v^{e}$ is $u^{e}$. However, given our choice of $j_{1}$ and $j_{2}$, none of the operations that happen on vertex sets between $V_{j_{1}}$ and $V_{j_{2}}$ involve $u^{e}$ or $v^{e}$, so if the operations taking place between $V_{j_{1}}$ and $V_{j_{2}}$ are legal in $\mathbb{V}$, they will remain legal in $\mathbb{V}^{\prime}$. ∎
Since we know that if we are dealing with a nontrivial sequence, there exists a solution $\mathbb{V}$ with at least one endpoint vertex removal (Observation 14), such that each endpoint vertex removal $u^{uv}$ is directly followed by an endpoint vertex addition $v^{uv}$ (Observations 15 and 16). Moreover, if $i$ is odd, vertex set $V_{i+1}$ in $\mathbb{V}$ is obtained from $V_{i}$ by removing an endpoint vertex $u^{uv}$, and vertex set $V_{i+2}$ is obtained from $V_{i+1}$ by adding the other endpoint vertex $v^{uv}$ associated with the same edge. It follows that that $|V_{i}|=|E|$ if $i$ is odd and $|V_{i+1}|=|E|-1$ if $i$ is even.

### v3

@@ 3.2 Hardness results
The instances of $k$-Compatible Ordering problem that we were efficiently able to solve so far (using Corollary 1) exhibited acyclicity within their input. In this section, we restrict the presence of cycles to represent cyclical constraints that naturally arise in the context of the RAMP problems. We prove below that $k$-Compatible Ordering is hard under each of those restrictions.
@@
Membership in $\mathsf{NP}$ follows from the fact that a labeled ordering $(\mathcal{S},\mathcal{L})$ serves as a certificate; verifying all constraints in $\mathcal{O}(n+m)$ time is straightforward (Observation 6).
@@
Let $c_{i}^{a}$ be a clause vertex labeled 1, and let $x_{j}$ be the associated variable. If $c_{i}^{a}$ is a positive clause vertex, given the arc $c_{i}^{a}x_{j}$ in $A_{1}$, $x_{j}$ must come before $c_{i}^{a}$ (Observation 2). Moreover, since we can now assume that $x_{j}$ comes before $c_{i}^{a}$, $x_{j}$ cannot be assigned label 2 due to the arc $x_{j}c^{a}_{i}$ in $A_{2}$. Otherwise, if $c_{i}^{a}$ is a negative clause vertex, given the arc $x_{j}c^{a}_{i}$ in $B_{1}$, $x_{j}$ must come after $c_{i}^{a}$ (Observation 2). Moreover, since we can now assume that $x_{j}$ comes after $c_{i}^{a}$, $x_{j}$ cannot be assigned label 1 due to the arc $c^{a}_{i}x_{j}$ in $B_{1}$. ∎
Let $C^{\ell=1}$ be the set of clause vertices that are assigned label 1. It follows from Observation 9 that $C^{\ell=1}$ cannot include negative and positive clause vertices with the same associated variable, as this would imply that the associated variable is included multiple times in the ordering, and is assigned two different labels. We partition $C^{\ell=1}$ into two sets of clause vertices, $C^{+,\ell=1}$ and $C^{-,\ell=1}$, based on polarity, i.e., $C^{+,\ell=1}$ (resp. $C^{-,\ell=1}$) is the set of positive (resp. negative) clause vertices in $C^{l=1}$. Let $X^{+,\ell=1}$ (resp. $X^{-,\ell=1}$) be the set of variable vertices associated with the clause vertices in $C^{+,\ell=1}$ (resp. $C^{-,\ell=1}$). Note that $|X^{+,\ell=1}|\leq|C^{+,\ell=1}|$ and $|X^{-,\ell=1}|\leq|C^{-,\ell=1}|$, since multiple clause vertices may be associated with the same variable vertex, and that $X^{+,\ell=1}\cap X^{-,\ell=1}=\emptyset$, which follows from Observation 9.
We observe that $X\setminus\left(X^{+,\ell=1}\bigcup X^{-,\ell=1}\right)$ need not be empty. We claim that setting each variable in $X^{+,\ell=1}$ to true and setting each variable in $X^{-,\ell=1}$ to false satisfies $\varphi$; the variables in $X\setminus\left(X^{+,\ell=1}\bigcup X^{-,\ell=1}\right)$ are don’t-care variables that can be set to either true or false. Suppose otherwise: there exists a clause $C_{i}$ in $\varphi$, that is not satisfied by the proposed truth value assignment. By Observation 4, a clause vertex $c_{i}^{a}$ of $C_{1}$, whose associated variable is $x_{j}$, is included in the ordering with label 1. Furthermore, by Observation 9, if $c^{a}_{i}$ is a positive (resp. negative) clause vertex, $x_{j}\in X^{+,\ell=1}$ (resp. $x_{j}\in X^{-,\ell=1}$), and is assigned the truth assignment true (resp. false) by our construction, which satisfies $c_{i}^{1}$, and therefore $C_{i}$, a contradiction. This concludes the proof of correctness of the reduction.
@@
The construction covered in Theorem 2 goes against the spirit of the problem, because it relies on the existence of directed digons in $B_{1}$ and directed triangles in $A_{2}$: directed cycles trivially prevent the existence of a solution where the cycle’s vertices are all assigned the same label (Observation 4). One might therefore wonder whether forbidding directed cycles within each of the directed graphs makes the problem easier to solve. Our next result shows that the answer is negative.
@@
    Arcs involving the global vertex $g$: For each dummy clause vertex $d_{i}$, add the arc $d_{i}g$ to $A_{2}$ and the arc $gd_{i}$ to $B_{2}$. For each dummy negative clause vertex $d_{i}^{a}$, add the arc $d_{i}^{a}g$ to $A_{1}$ and the arc $gd_{i}^{a}$ to $B_{1}$. These arcs guarantee dummy vertices obtain particular labels, as explained in Observation 5, and prevent cycles involving $g$.
@@
($\Leftarrow$) The proof for the backward direction of Theorem 2 relied on the fact that one of $c_{i}^{1},c_{i}^{2}$ and $c_{i}^{3}$ is assigned label 1 for each clause $C_{i}$, as well as Observation 9. We show that the two observations still hold under the new construction, which is sufficient to prove the existence of a satisfying assignment for $\varphi$.
To prove that, for each clause $C_{i}$, one of $c_{i}^{1},c_{i}^{2},c_{i}^{3}$ is assigned label 1 under the new construction, suppose, for the sake of contradiction, that there exists a clause $C_{i}$ whose clause vertices are all assigned label 2 in the ordering, which implies that its non-dummy clause vertices are in the following relative order: $c_{i}^{3},c_{i}^{2},c_{i}^{1}$. Since $d_{i}$ is assigned label 1 (Observation 5), it will have to come after $c_{i}^{1}$, and therefore, after $c_{i}^{3}$, which violates a constraint on $c_{i}^{3}$ because of the arc $c_{i}^{3}d_{i}$ in $A_{1}$ (Observation 2). As for Observation 9, the previous proof is still applicable in the new construction, because the proof relies on the arcs constructed between variable vertices and non-dummy clause vertices, which were not modified.
@@
In the construction of Theorem 3, cycles no longer exist in individual directed graphs, but rather, in the union of the directed graphs. By Corollary 1, we already know that the problem admits a polynomial-time solution when the union of $\mathcal{G}$ is acyclic (and actually even when there exists a single $i$ for which $A_{i}\cup B_{i}$ is acyclic). So the natural next step would be looking for intermediate instances. Next, we prove that the problem remains $\mathsf{NP}$-complete even on instances whose union is acyclic after a simple transformation of the instance. Given a directed graph $G=(V,E)$, we define the reversed graph of $G$ as the graph $(V,\{v_{j}v_{i}$ $|$ $v_{i}v_{j}\in E\})$: we will reverse the $B$ graphs.
@@
($\Leftarrow$) Let $(\mathcal{S},\mathcal{L})$ be the valid labeled ordering that solves the constructed $3$-Compatible Ordering instance. In addition to the vertex $t_{1}$ coming before the vertex $t_{2}$ in $\mathcal{S}$, since the arc $t_{2}t_{1}$ exists in all $A$ graphs (Observation 3), $\mathcal{S}$ has the following characteristics:
Observation 10.
@@
Observation 10 is due to $t_{1}$ coming before $t_{2}$, and the arcs $C_{i}t_{1}$ in the $A$ graphs, $t_{2}C_{i}$ in the $B$ graphs for each clause vertex $C_{i}$. As for Observation 11, for each variable vertex $x_{j}$, the arc $t_{1}x_{j}$ in $B_{1}$ means that $x_{j}$ must come before $t_{1}$ if $x_{j}$ is labeled 1 (Observation 2), the arcs $x_{j}t_{1}$ and $t_{1}x_{j}$ in $A_{2}$ and $B_{2}$ respectively mean that $x_{j}$ is never labeled 2 (Observation 5), and the arc $x_{j}t_{2}$ in $A_{3}$ means that $x_{j}$ must come after $t_{2}$ if $x_{j}$ is labeled 3 (Observation 2).
In order to derive a satisfying assignment, we focus on the clause vertices, as well as the labels that they are assigned. Let $\ell_{1},\ldots,\ell_{p}$ be the labels assigned to $C_{1},\ldots,C_{p}$, respectively. Consider the set of literals $C^{sat}=\{c^{\ell_{1}}_{1},\ldots,c^{\ell_{p}}_{p}\}$. We will partition those literals into two sets $C^{T}$ and $C^{F}$, based on whether the arc between $c_{i}^{\ell_{i}}$ and its associated variable, in one direction or the other, is in an $A$ graph or a $B$ graph, respectively. We then define $X^{T}$ (resp. $X^{F}$) as the set of variables in $\varphi$ with at least one literal in $C^{T}$ (resp. $C^{F}$). We observe that a variable vertex $x$ cannot belong to both $X^{T}$ and $X^{F}$, i.e., $X^{T}\cap X^{F}=\emptyset$, because an arc from a clause vertex in $C^{sat}$ to $x$ in an $A$ graph implies that $x$ comes before $t_{1}$ in $\mathcal{S}$, and an arc from the same clause vertex in $C^{sat}$ to $x$ in a $B$ graph implies that $x$ comes after $t_{2}$ in $\mathcal{S}$ (Observations 211), a contradiction, since we already established that $t_{1}$ comes before $t_{2}$ in every valid labeled ordering.
@@ 5 Finding $k$-compatible set arrangements
The proof of Observation 13 is identical to that of Observation 12 because vertex additions and vertex removals are treated identically as far as constraint satisfaction is concerned, i.e., the removal of a vertex $v$ from $V_{i}$ is legal if and only if the addition of a vertex $v$ to $V_{i}\setminus\{v\}$ is legal.
@@
Analyzing whether the vertex removal we propose is legal will omit any mention of the $A$ graphs, because we have already established that a vertex removal cannot violate the removed vertex’s sink constraint (Observation 12):
@@
If $V_{\text{s}}=V_{\text{t}}$, there is nothing to do. Otherwise, we can refer to Observation 13 to conclude the existence of at least one vertex removal in $\mathbb{V}$. ∎
Observation 16.
@@
Suppose otherwise, for the sake of contradiction. If the removal of $v^{e}$ is the last addition/removal operation involving any of $v^{e}$ and $u^{e}$ in $\mathbb{V}$, we can find a shorter sequence of vertex sets that is a solution to the instance. If $\mathbb{V}=V_{1},\ldots,V_{j_{1}},V_{j_{1}+1},V_{j_{1}+2},\ldots,V_{q}$ is a solution to the instance, so is $\mathbb{V}^{\prime}=V_{1},\ldots,V_{j_{1}},V_{j_{1}+2}\cup\{v^{e}\},\ldots,V_{q}\cup\{v^{e}\}$. Undoing the removal of $v^{e}$ can only cause sink constraint violations for the vertices that are subsequently added or removed (Observation 14), and the only vertex whose sink constraint involves $v^{e}$ is $u^{e}$. A similar argument can be used to shorten the vertex set sequence that is a solution to the instance if the removal of $v^{e}$ is followed by the addition of $v^{e}$ to some vertex set $V_{j_{3}}$, where $j_{3}>j_{1}$. ∎
In what follows, we assume that $\mathbb{V}$ is a nontrivial shortest solution whose structure adheres to that of the solution in the statement of Observation 16.
Observation 17.
@@
We modify $\mathbb{V}$ to obtain $\mathbb{V}^{\prime}$ by moving $v^{e}$ directly after the removal operation of $u^{e}$, which can be thought of as the undoing of a vertex addition of $v^{e}$, followed by a vertex addition of $v^{e}$, both to a sequence of vertex sets. This modification can only violate sink constraints (Observation 14), and the only vertex that is potentially affected by the earlier inclusion of $v^{e}$ is $u^{e}$. However, given our choice of $j_{1}$ and $j_{2}$, none of the operations that happen on vertex sets between $V_{j_{1}}$ and $V_{j_{2}}$ involve $u^{e}$ or $v^{e}$, so if the operations taking place between $V_{j_{1}}$ and $V_{j_{2}}$ are legal in $\mathbb{V}$, they will remain legal in $\mathbb{V}^{\prime}$. ∎
Since we know that if we are dealing with a nontrivial sequence, there exists a solution $\mathbb{V}$ with at least one endpoint vertex removal (Observation 15), such that each endpoint vertex removal $u^{uv}$ is directly followed by an endpoint vertex addition $v^{uv}$ (Observations 16 and 17). Moreover, if $i$ is odd, vertex set $V_{i+1}$ in $\mathbb{V}$ is obtained from $V_{i}$ by removing an endpoint vertex $u^{uv}$, and vertex set $V_{i+2}$ is obtained from $V_{i+1}$ by adding the other endpoint vertex $v^{uv}$ associated with the same edge. It follows that that $|V_{i}|=|E|$ if $i$ is odd and $|V_{i+1}|=|E|-1$ if $i$ is even.